OpenStudy (anonymous):

Petrolyn motor oil is a combination of natural oil and synthetic oil. It contains 7 liters of natural oil for every 4 liters of synthetic oil. If 399 liters of natural oil are used, how many liters of Petrolyn oil will be made?

OpenStudy (welshfella):

the fraction of natural oil in the petrolyn = 7/11 fraction of synthetic oil = 4/11